# Mitigating Visual Knowledge Forgetting in MLLM Instruction-tuning via Modality-decoupled Gradient Descent

### v1

@@ Abstract.
Recent MLLMs have shown emerging visual understanding and reasoning abilities after being pre-trained on large-scale multimodal datasets. Unlike pre-training, where MLLMs receive rich visual-text alignment, instruction-tuning is often text-driven with weaker visual supervision, leading to the degradation of pre-trained visual understanding and causing visual forgetting. Existing approaches, such as direct fine-tuning and continual learning methods, fail to explicitly address this issue, often compressing visual representations and prioritizing task alignment over visual retention, which further worsens visual forgetting. To overcome this limitation, we introduce a novel perspective leveraging effective rank to quantify the degradation of visual representation richness, interpreting this degradation through the information bottleneck principle as excessive compression that leads to the degradation of crucial pre-trained visual knowledge. Building on this view, we propose a modality-decoupled gradient descent (MDGD) method that regulates gradient updates to maintain the effective rank of visual representations while mitigating the over-compression effects described by the information bottleneck. By explicitly disentangling the optimization of visual understanding from task-specific alignment, MDGD preserves pre-trained visual knowledge while enabling efficient task adaptation. To enable lightweight instruction-tuning, we further develop a memory-efficient fine-tuning approach using gradient masking, which selectively updates a subset of model parameters to enable parameter-efficient fine-tuning (PEFT), reducing computational overhead while preserving rich visual representations. Extensive experiments across various downstream tasks and backbone MLLMs demonstrate that MDGD effectively mitigates visual forgetting from pre-trained tasks while enabling strong adaptation to new tasks.
Visual Knowledge Forgetting, Multimodal Large Language Models
1. Introduction
Multimodal large language models (MLLMs) enhanced visual understanding and reasoning by pre-training on large-scale multimodal datasets with comprehensive visual descriptions that integrats textual and visual knowledge (Liu et al., 2024b; Yao et al., 2024; Li et al., 2023b; Bai et al., 2023; Liu et al., 2024d). These models achieve strong performance across various vision-language tasks, such as visual question answering (Jin et al., 2024), multimodal reasoning (Zhang et al., 2024; Jiang et al., 2024b; Yan et al., 2024), multimodal recognition (Shenoy et al., 2024; Wu et al., 2024e), personalized multimodality (Wu et al., 2024c), and document intelligence (Jin et al., 2024; Shenoy et al., 2024). However, adapting pre-trained MLLMs to downstream tasks via instruction-tuning (Wu et al., 2024a; Li et al., 2024, 2023c; Panagopoulou et al., 2023; Liu et al., 2024d) presents a critical challenge of visual forgetting. Unlike pre-training, where models receive rich visual-text alignment, instruction-tuning is often text-driven with limited direct visual supervision. This shift in training focus leads to the degradation of pre-trained visual encoding (Zhou et al., 2024; Niu et al., 2024; Wu et al., 2024a; Ko et al., 2023), negatively impacting model generalizability across downstream tasks that require strong visual knowledge (Bai et al., 2024; Huang et al., 2024). Addressing this challenge is essential for ensuring MLLMs retain their visual capabilities while aligning with new tasks efficiently.
While several approaches have attempted to mitigate catastrophic forgetting in neural networks through direct fine-tuning and continual learning methods (Shi et al., 2024; Wu et al., 2024b; Zhu et al., 2024; Zheng et al., 2024), these methods often overlook the unique challenge of preserving visual knowledge in multimodal large language models (MLLMs). Directly fine-tuning MLLMs on new tasks often leads to overfitting to textual instructions while inadvertently suppressing visual representations (Zhai et al., 2023). Existing continual learning strategies such as regularization and replay methods tend to focus on retaining language-based knowledge, neglecting the critical trade-off between compressing visual representations and aligning them with task-specific instructions (Zhou et al., 2024; Niu et al., 2024; Wu et al., 2024a; Ko et al., 2023), leading to the degradation of pre-trained visual knowledge. Task-orthogonal gradient descent techniques have shown promise in disentangling gradients for multi-task optimization. However, their practical application in MLLMs poses unique challenges. MLLMs are pre-trained on vast and heterogeneous multimodal datasets (Liu et al., 2024b; Li et al., 2023b; Bai et al., 2023), where it is challenging to precisely isolate task-specific gradients, causing the components critical for visual understanding to become entangled with other features.
@@ 1. Introduction
Figure 1. The top-10 image tokens with the highest effective ranks on OKVQA and POPE encoded by LLaVA, and PathVQA and POPE encoded by MiniCPM. We compare pretrained, finetuned, and MDGD-finetuned models. Effective rank (Wei et al., 2024) quantifies representation richness, and we novelly use it to analyze visual degradation in instruction-tuned MLLMs. Results show that MDGD preserves higher effective rank, mitigating visual forgetting.
@@
To gain a fundamental view of the challenge of visual knowledge forgetting in MLLM instruction tuning, we adopt an information bottleneck (IB) perspective that characterizes the trade-off between retaining input information and ensuring output predictiveness (Tishby et al., 2000). To investigate the degradation of crucial pre-trained visual knowledge, we introduce a novel perspective that leverages effective rank to quantify the richness of the encoded visual representation from MLLMs. Specifically, we illustrate the visual forgetting problem in Figure 1, where we observe a consistent effective rank reduction problem caused by MLLM instruction tuning. Based on this view, we propose a modality-decoupled gradient descent (MDGD) method, which disentangles the optimization of visual understanding from task-specific alignment, MDGD regulates gradient updates to maintain the effective rank of visual representations compared with pre-trained MLLMs, while mitigating the over-compression effects described by the information bottleneck. Intuitively, visual forgetting occurs due to the shift from rich multimodal pre-training to instruction-tuning, where text-based supervision dominates without direct visual supervision. By explicitly decoupling the task-specific alignment with visual representation learning, MDGD preserves expressive and robust visual features. To further improve efficiency in instruction-tuning, we introduce a memory-efficient fine-tuning strategy using gradient masking, which selectively updates a subset of model parameters for parameter-efficient fine-tuning (PEFT). This approach reduces computational overhead while ensuring that crucial pre-trained visual representations are retained.
@@
2. Related Work
2.1. Visual Knowledge Forgetting in MLLMs
Catastrophic forgetting, a persistent challenge in continual learning, occurs when a model forgets previous knowledge while learning new tasks, reducing its performance on earlier tasks (Wang et al., 2023). This issue has gained attention in LLMs due to the growing need for continual pre-training and instruction tuning (Wu et al., 2024b; Luo et al., 2023). MLLMs, which integrate multiple modalities through feature encoders projecting inputs into the LLM’s token space, are also prone to catastrophic forgetting (Zhai et al., 2023). While several methods have attempted to address this issue by adapting continual learning techniques, ranging from fine-tuning and task-orthogonal gradient descent to knowledge distillation and replay-based strategies (Shi et al., 2024; Wu et al., 2024b; Zhu et al., 2024; Zheng et al., 2024), these approaches often fall short of preserving rich visual representations. For instance, fine-tuning MLLMs on new tasks tends to overfit textual instructions, inadvertently suppressing visual features, and even parameter-efficient adaptations like LoRA have been shown to suffer from forgetting (Fawi, 2024; Liu et al., 2024c). Model Tailor (Zhu et al., 2024) addresses forgetting by adapting the LLM backbone across reasoning tasks but neglects the critical visual knowledge forgetting problem, which may lead to visual hallucination or deficiency problems while generalizing to various tasks (Zhai et al., 2023). In contrast, our method offers a more principled and synchronized approach to instruction tuning that jointly optimizes the alignment between the visual encoder and the LLM, effectively preserving pre-trained visual knowledge while mitigating the degradation of visual representation learning in previous works.
2.2. Information Theory in LLMs
The Information Bottleneck (IB) method (Tishby et al., 2000) in Large Language Models (LLMs) focuses on compressing input data while preserving information relevant to the target output (Delétang et al., 2023; Valmeekam et al., 2023; Wei et al., 2024; Wu et al., 2022). Previous works employ IB to extract robust task-specific features (Zhang et al., 2022; Wu et al., 2024d) and enable feature attribution (Li et al., 2022; Jiang et al., 2020) However such methods focus on information compression in language models (Yang et al., 2025), while the unique visual knowledge forgetting problem cannot be directly viewed and interpreted. In addition, existing information-theoretic transfer learning methods (Tseng et al., 2024; Wu et al., 2024d; Ling et al., 2024) cannot be directly applied to MLLMs, where the multimodal knowledge is entangled and fused by a dominant LLM. In contrast to these prior works, our approach leverages effective rank to quantify and mitigate the over-compression of visual representations in MLLMs. Specifically, our modality-decoupled gradient descent (MDGD) method explicitly decouples visual optimization from task-specific alignment, which cannot be enabled by existing IB-based methods for LLMs.
3. Preliminary
@@
(2) $\mathcal{L}_{vl}(\theta)=-\sum_{t=1}^{T}\log\pi_{\theta}\left(y_{t}\mid X^{v},X^{l},y_{<t}\right),$
@@
An Information Bottleneck Perspective on Visual Knowledge Forgetting. In multimodal models, the information bottleneck (Mai et al., 2022) (IB) framework provides a powerful lens to understand how representations are formed. In our setting, the IB principle seeks a representation $Z$ that is maximally informative about the output $y$ while discarding irrelevant details from the inputs. For an MLLM that processes visual inputs $X^{v}$ and textual inputs $X^{l}$, a full IB objective might take the form:
(3) $\min_{\theta}\quad\mathcal{L}_{\text{IB}}^{\text{vision}}(\theta)=-I(y;Z)+\beta\,I(X^{v};Z).$
@@
where $I(\cdot;\cdot)$ denotes mutual information and $\beta$ controls the trade-off between predictive power and compression. This formulation explicitly highlights the risk of discarding visual details when the model is optimized primarily to predict $y$.
@@
Effective Rank as a Measure of Representation Richness. To quantify the information content retained in a representation, we use the effective rank metric (Roy and Vetterli, 2007). Given a representation matrix $Z$ whose singular values are $\{\sigma_{i}\}$, the effective rank is defined as:
(4) $\text{erank}(Z)=\exp\Biggl{(}-\sum_{i}p_{i}\log p_{i}\Biggr{)},\quad\text{with}\quad p_{i}=\frac{\sigma_{i}}{\sum_{j}\sigma_{j}}.$
@@
This measure, based on the entropy of the singular value distribution, captures the “richness” or intrinsic dimensionality of $Z$. A higher effective rank indicates that the representation spans a larger subspace, whereas a lower effective rank implies that the representation has been overly compressed.
4. Visual Forgetting in MLLM Instruction-tuning
$Z\sim p(\cdot\mid X^{v},X^{l}),$
During instruction tuning, the model is optimized primarily to predict the target $y$. As described in Eq. (3), the IB objective introduces a trade-off between retaining visual information $I(X^{v};Z)$ and ensuring that $Z$ remains predictive of $y$ via $I(y;Z)$ (Jiang et al., 2024a). In practice, however, instruction-tuning datasets are predominantly text-driven; thus, the learned visual representation $Z^{v}$ receives only indirect and often weaker supervision (Wang et al., 2024).
$\mathcal{Z}_{\theta}^{v}=\left\{Z^{v}_{\theta}:Z\sim p_{\theta}(\cdot\mid X^{v},X^{l}),\quad(X^{v},X^{l})\in D\right\},$
(5) $\text{erank}(\mathcal{Z}_{\theta}^{v})<\text{erank}(\mathcal{Z}_{0}^{v}).$
5. MDGD: Modality-Decoupled Gradient Regularization and Descent
@@
Motivated by the visual forgetting problem caused by the degradation of multimodal encoding in Eq. (5), we introduce a modality-decoupling gradient regularization (MDGD) to approximate orthogonal gradients between visual understanding drift and downstream task optimization. Specifically, leveraging modality-decoupled gradients $\bar{g}_{\theta}$ and $\bar{g}_{\phi}$ derived from the current MLLM and a pre-trained MLLM respectively, we propose a gradient regularization term $\tilde{g}_{\theta}$ for more efficient multimodal instruction tuning, which promotes the alignment of downstream tasks while mitigating visual forgetting (Zhu et al., 2024). Since MDGD requires the estimation of parameter gradients, we could not directly apply parameter-efficient fine-tuning methods (e.g., LoRA (Hu et al., 2021)). Thus, we alternatively formulate the regularization as a gradient mask $M_{\tilde{g}_{\theta}}$, which allows efficient fine-tuning only on a subset of masked model parameters.
5.1. Modality Decoupling
@@
Based on the information bottleneck objective in Eq. (3), the objective encourages the model to maximize $I(y;Z)$ while compressing $I(X^{v};Z)$ (Tishby et al., 2000; Alemi et al., 2016). In practice, this compression may discard useful visual details, leading to visual forgetting. To mitigate such compression and preserve the pre-trained visual knowledge, we follow the KL divergence loss $D_{\text{KL}}\Bigl{(}\mu_{\phi}(X^{v})\,\Big{\|}\,\pi_{\theta}(X^{v})\Bigr{)}$ to constrain the current model’s visual representation $\pi_{\theta}(X^{v})$ to remain close to the pre-trained distribution $\mu_{\phi}(X^{v})$, thereby preserving the mutual information $I(X^{v};Z)$ that would otherwise be reduced by the compression (Hinton, 2015; Lopez et al., 2018). However, since MLLMs cannot directly track the distributions of image tokens, we instead introduce an auxiliary loss function
(6) $\mathcal{L}_{v}(\phi,\theta)=\|\mu(X^{v}|\phi)-\pi(X^{v}|\theta)\|_{1},$
@@
which approximates the KL divergence loss (Zhu et al., 2022a, 2017) by penalizing discrepancies between the pre-trained visual representation and that obtained during instruction tuning.
@@
(7) $\bar{g}_{\phi}$ $=\nabla_{\phi}\mathcal{L}_{vl}(\phi)-\frac{\nabla_{\phi}\mathcal{L}_{vl}(\phi)^{\top}h_{\phi}}{\|h_{\phi}\|^{2}}\cdot h_{\phi},$
(8) $\bar{g}_{\theta}$ $=\nabla_{\theta}\mathcal{L}_{vl}(\theta)-\frac{\nabla_{\theta}\mathcal{L}_{vl}(\theta)^{\top}h_{\theta}}{\|h_{\theta}\|^{2}}\cdot h_{\theta},$
@@
Figure 2. Illustration of the proposed method. To mitigate suboptimal optimization and prevent visual forgetting, we first project $\nabla_{\theta}\mathcal{L}_{vl}$ onto the direction orthogonal to $\nabla_{\theta}\mathcal{L}_{v}$, obtaining $\bar{g}_{\theta}$. Next, we project $\bar{g}_{\theta}$ onto the direction of $\bar{g}_{\phi}$, yielding $\tilde{g}_{\theta}$. This process guides the gradient towards the optimal region without visual forgetting.
@@
5.2. Regularized Gradient Descent
The auxiliary loss in Eq. (6) preserves the visual representation at a distribution level via the feature alignment auxiliary loss in Eq. (6). However, the information bottleneck framework indicates that the gradient component compressing $I(X^{v};Z)$ (i.e., $\nabla_{\theta}I(X^{v};Z)$), can harm visual preservation by reducing the effective rank of the features (Achille and Soatto, 2018; Lee et al., 2021).
To address this compression-induced drift, we incorporate an orthogonal gradient as a regularize. Motivated by multi-task orthogonal gradient optimization (Yu et al., 2020; Zhu et al., 2022b; Dong et al., 2022), we leverage the gradient $\bar{g}_{\phi}$ from the pre-trained model $\mu_{\phi}$, which reflects the accumulated visual drift and approximates a global orthogonal learning effect in the downstream task. We then project the current model’s gradient onto this direction:
(9) $\tilde{g}_{\theta}=\frac{\bar{g}_{\theta}^{\top}\bar{g}_{\phi}}{\|\bar{g}_{\phi}\|^{2}}\cdot\bar{g}_{\phi}.$

### v2

@@ Abstract
Recent MLLMs have demonstrated strong visual understanding and reasoning after large-scale multimodal pre-training. However, instruction-tuning is typically text-driven with limited visual supervision, leading to significant visual forgetting and degradation of pre-trained visual knowledge. Existing fine-tuning and continual learning methods compress visual representations and emphasize task alignment over visual retention, failing to address this challenge. We present a novel perspective using effective rank to quantify the loss of visual representation richness, framing visual forgetting as excessive compression under the information bottleneck principle. To address this, we propose modality-decoupled gradient descent (MDGD), which regulates gradient updates to preserve the effective rank of visual features and explicitly disentangles visual learning from task-specific alignment. We further introduce a memory-efficient fine-tuning variant using gradient masking for parameter-efficient adaptation. Extensive experiments show that MDGD effectively mitigates visual forgetting across downstream tasks and models, maintaining pre-trained visual knowledge while supporting strong task adaptation.
Mitigating Visual Knowledge Forgetting in MLLM Instruction-tuning via Modality-decoupled Gradient Descent
Junda Wu1, Yuxin Xiong1, Xintong Li1, Yu Xia1, Ruoyu Wang2, Yu Wang1 Tong Yu3, Sungchul Kim3, Ryan Rossi3, Lina Yao2, Jingbo Shang1, Julian McAuley1 1UC San Diego  2University of New South Wales  3Adobe Research {juw069,y7xiong,xil240,yux078,yuw164,jshang,jmcauley}@ucsd.edu {ruoyu.wang5,lina.yao}@unsw.edu.au{tyu,sukim,ryrossi}@adobe.com
1 Introduction
@@
Figure 1: The top-10 image tokens with the highest effective ranks on OKVQA and POPE encoded by LLaVA, and PathVQA and POPE encoded by MiniCPM. We compare pretrained, finetuned, and MDGD-finetuned models. Effective rank Wei et al. (2024) quantifies representation richness, which show that MDGD preserves higher effective rank, mitigating visual forgetting.
Multimodal large language models (MLLMs) enhanced visual understanding and reasoning by pre-training on large-scale multimodal datasets with comprehensive visual descriptions that integrate textual and visual knowledge Liu et al. (2024b); Yao et al. (2024); Li et al. (2023b); Bai et al. (2023); Liu et al. (2024d); Wu et al. (2024d). These models achieve strong performance across various vision-language tasks, such as visual question answering Jin et al. (2024); Wu et al. (2025b), multimodal reasoning Zhang et al. (2024); Jiang et al. (2024b); Yan et al. (2024), multimodal recognition Shenoy et al. (2024); Wu et al. (2024d), and personalized multimodality Wu et al. (2024b, 2025a); Huang et al. (2025). However, adapting pre-trained MLLMs to downstream tasks via instruction-tuning Wu et al. (2024a); Li et al. (2024, 2023a); Panagopoulou et al. (2023); Liu et al. (2024d) presents a critical challenge of visual forgetting. Unlike pre-training, where models receive rich visual-text alignment, instruction-tuning is often text-driven with limited direct visual supervision. This shift in training focus leads to the degradation of pre-trained visual encoding Zhou et al. (2024); Niu et al. (2024); Wu et al. (2024a); Ko et al. (2023), negatively impacting model generalizability across downstream tasks that require strong visual knowledge Bai et al. (2024); Huang et al. (2024). Addressing this challenge is essential for ensuring MLLMs retain their visual capabilities while aligning with new tasks efficiently.
While several approaches have attempted to mitigate catastrophic forgetting in neural networks through direct fine-tuning and continual learning methods Shi et al. (2024); Wu et al. (2024e); Zhu et al. (2024); Zheng et al. (2024), these methods often overlook the unique challenge of preserving visual knowledge in multimodal large language models (MLLMs). Directly fine-tuning MLLMs on new tasks often leads to overfitting to textual instructions while inadvertently suppressing visual representations (Zhai et al., 2023). Existing continual learning strategies, such as regularization and replay methods, tend to focus on retaining language-based knowledge, neglecting the trade-off between compressing visual representations and aligning them with task-specific instructions Zhou et al. (2024); Niu et al. (2024); Wu et al. (2024a); Ko et al. (2023), leading to the degradation of pre-trained visual knowledge. Task-orthogonal gradient descent techniques have shown promise in disentangling gradients for multi-task optimization. However, their practical application in MLLMs poses unique challenges. MLLMs are pre-trained on vast and heterogeneous multimodal datasets Liu et al. (2024b); Li et al. (2023b); Bai et al. (2023), where it is challenging to isolate task-specific gradients, causing the components critical for visual understanding to become entangled with other features.
@@
To gain a fundamental view of the challenge of visual knowledge forgetting in MLLM instruction tuning, we adopt an information bottleneck (IB) perspective that characterizes the trade-off between retaining input information and ensuring output predictiveness Tishby et al. (2000). To investigate the degradation of crucial pre-trained visual knowledge, we introduce a novel perspective that leverages effective rank to quantify the richness of the encoded visual representation from MLLMs. Specifically, we illustrate the visual forgetting problem in Figure 1, where we observe a consistent effective rank reduction problem caused by MLLM instruction tuning. Based on this view, we propose a modality-decoupled gradient descent (MDGD) method, which disentangles the optimization of visual understanding from task-specific alignment, MDGD regulates gradient updates to maintain the effective rank of visual representations compared with pre-trained MLLMs, while mitigating the over-compression effects described by the information bottleneck. Intuitively, visual forgetting occurs due to the shift from rich multimodal pre-training to instruction-tuning, where text-based supervision dominates without direct visual supervision. By explicitly decoupling the task-specific alignment with visual representation learning, MDGD preserves expressive and robust visual features. To further improve efficiency in instruction-tuning, we introduce a memory-efficient fine-tuning strategy using gradient masking, which selectively updates a subset of model parameters for parameter-efficient fine-tuning (PEFT). This approach reduces computational overhead while ensuring that crucial pre-trained visual representations are retained.
@@
1.1 Visual Knowledge Forgetting in MLLMs
Catastrophic forgetting, where a model loses previous knowledge while learning new tasks, is a major challenge in continual learning Wang et al. (2023). This problem is now widely recognized in LLMs and MLLMs Wu et al. (2024e); Luo et al. (2023); Zhai et al. (2023). Although various methods, including fine-tuning, task-orthogonal gradient descent, knowledge distillation, and replay, have been adapted to mitigate forgetting Shi et al. (2024); Wu et al. (2024e); Zhu et al. (2024); Zheng et al. (2024), they often fail to preserve rich visual features. Fine-tuning on new tasks tends to overfit text and suppress visual information, while parameter-efficient methods like LoRA also suffer from forgetting Fawi (2024); Liu et al. (2024c). Model Tailor Zhu et al. (2024) adapts the LLM backbone but does not address visual knowledge forgetting, which can lead to hallucination or degraded generalization Zhai et al. (2023). In contrast, our approach synchronizes the training of the visual encoder and LLM, preserving pre-trained visual knowledge during instruction tuning.
1.2 Information Theory in LLMs
The Information Bottleneck (IB) principle Tishby et al. (2000) has been used in LLMs to compress input while retaining task-relevant information Delétang et al. (2023); Valmeekam et al. (2023); Wei et al. (2024); Wu et al. (2022). Prior works use IB to extract robust features Zhang et al. (2022); Wu et al. (2024c) and enable feature attribution Li et al. (2022); Jiang et al. (2020), but these focus on language models, not multimodal settings Yang et al. (2025). Existing information-theoretic transfer learning methods Tseng et al. (2024); Wu et al. (2024c); Ling et al. (2024) also do not address the unique challenges of MLLMs, where modalities are deeply entangled. Our method instead uses effective rank to measure and counteract visual representation compression. The proposed MDGD explicitly decouples visual learning from task alignment, going beyond previous IB-based approaches.
2 Preliminary
@@
$\mathcal{L}_{vl}(\theta)=-\sum_{t=1}^{T}\log\pi_{\theta}\left(y_{t}\mid X^{v},X^{l},y_{<t}\right),$ (2)
@@
An Information Bottleneck Perspective on Visual Knowledge Forgetting. In multimodal models, the information bottleneck Mai et al. (2022) (IB) framework provides a powerful lens to understand how representations are formed. In our setting, the IB principle seeks a representation $Z$ that is maximally informative about the output $y$ while discarding irrelevant details from the inputs. For an MLLM that processes visual inputs $X^{v}$ and textual inputs $X^{l}$, a full IB objective might take the form:
$\min_{\theta}\quad\mathcal{L}_{\text{IB}}^{\text{vision}}(\theta)=-I(y;Z)+\beta\,I(X^{v};Z).$ (3)
@@
where $I(\cdot;\cdot)$ denotes mutual information and $\beta$ controls the trade-off between predictive power and compression. This formulation explicitly highlights the risk of discarding visual details when the model is optimized primarily to predict $y$. Based on the information bottleneck view, we further analyze the visual forgetting problem in Appendix C.
@@
Effective Rank as a Measure of Representation Richness. To quantify the information content retained in a representation, we use the effective rank metric Roy and Vetterli (2007). Given a representation matrix $Z$ whose singular values are $\{\sigma_{i}\}$, the effective rank is defined as:
$\text{erank}(Z)=\exp\Biggl(-\sum_{i}p_{i}\log p_{i}\Biggr),$ (4)
@@
where $p_{i}=\sigma_{i}/\sum_{j}\sigma_{j}$. This measure, based on the entropy of the singular value distribution, captures the “richness” or intrinsic dimensionality of $Z$. A higher effective rank indicates that the representation spans a larger subspace, whereas a lower effective rank implies that the representation has been overly compressed.
3 MDGD: Modality-Decoupled Gradient Regularization and Descent
@@
Motivated by the visual forgetting problem caused by the degradation of multimodal encoding in Eq. (12), we introduce a modality-decoupling gradient regularization (MDGD) to approximate orthogonal gradients between visual understanding drift and downstream task optimization. Specifically, leveraging modality-decoupled gradients $\bar{g}_{\theta}$ and $\bar{g}_{\phi}$ derived from the current MLLM and a pre-trained MLLM respectively, we propose a gradient regularization term $\tilde{g}_{\theta}$ for more efficient multimodal instruction tuning, which promotes the alignment of downstream tasks while mitigating visual forgetting Zhu et al. (2024). Since MDGD requires the estimation of parameter gradients, we could not directly apply parameter-efficient fine-tuning methods (e.g., LoRA Hu et al. (2021)). Thus, we alternatively formulate the regularization as a gradient mask $M_{\tilde{g}_{\theta}}$, which allows efficient fine-tuning only on a subset of masked model parameters.
3.1 Modality Decoupling
@@
Based on the information bottleneck objective in Eq. (3), the objective encourages the model to maximize $I(y;Z)$ while compressing $I(X^{v};Z)$ Tishby et al. (2000); Alemi et al. (2016). In practice, this compression may discard useful visual details, leading to visual forgetting. To mitigate such compression and preserve the pre-trained visual knowledge, we follow the KL divergence loss $D_{\text{KL}}\Bigl(\mu_{\phi}(X^{v})\,\Big\|\,\pi_{\theta}(X^{v})\Bigr)$ to constrain the current model’s visual representation $\pi_{\theta}(X^{v})$ to remain close to the pre-trained distribution $\mu_{\phi}(X^{v})$, thereby preserving the mutual information $I(X^{v};Z)$ that would otherwise be reduced by the compression Hinton (2015); Lopez et al. (2018). However, since MLLMs cannot directly track the distributions of image tokens, we instead introduce an auxiliary loss function
$\mathcal{L}_{v}(\phi,\theta)=\|\mu(X^{v}|\phi)-\pi(X^{v}|\theta)\|_{1},$ (5)
@@
which approximates the KL divergence loss Zhu et al. (2022b, 2017) by penalizing discrepancies between the pre-trained visual representation and that obtained during instruction tuning.
@@
$\bar{g}_{\phi}$ $=\nabla_{\phi}\mathcal{L}_{vl}(\phi)-\frac{\nabla_{\phi}\mathcal{L}_{vl}(\phi)^{\top}h_{\phi}}{\|h_{\phi}\|^{2}}\cdot h_{\phi},$ (6)
$\bar{g}_{\theta}$ $=\nabla_{\theta}\mathcal{L}_{vl}(\theta)-\frac{\nabla_{\theta}\mathcal{L}_{vl}(\theta)^{\top}h_{\theta}}{\|h_{\theta}\|^{2}}\cdot h_{\theta},$ (7)
@@
3.2 Regularized Gradient Descent
The auxiliary loss in Eq. (5) preserves the visual representation at a distribution level via the feature alignment auxiliary loss in Eq. (5). However, the information bottleneck framework indicates that the gradient component compressing $I(X^{v};Z)$ (i.e., $\nabla_{\theta}I(X^{v};Z)$), can harm visual preservation by reducing the effective rank of the features Achille and Soatto (2018); Lee et al. (2021).
To address this compression-induced drift, we incorporate an orthogonal gradient as a regularize. Motivated by multi-task orthogonal gradient optimization Yu et al. (2020); Zhu et al. (2022a); Dong et al. (2022), we leverage the gradient $\bar{g}_{\phi}$ from the pre-trained model $\mu_{\phi}$, which reflects the accumulated visual drift and approximates a global orthogonal learning effect in the downstream task. We then project the current model’s gradient onto this direction:
$\tilde{g}_{\theta}=\frac{\bar{g}_{\theta}^{\top}\bar{g}_{\phi}}{\|\bar{g}_{\phi}\|^{2}}\cdot\bar{g}_{\phi}.$ (8)